OpenStudy (anonymous):

How do you solve t=5π

OpenStudy (anonymous):

are you solving for t??? if so the solution is \[5\pi \]

OpenStudy (anonymous):

haha no, it says find the terminal point p(x,y) on the unit circle determined by each real number t.

OpenStudy (jdoe0001):

what's the terminal point for say \(\bf 2\pi \quad ?\)

OpenStudy (anonymous):

(1,0)

OpenStudy (anonymous):

oh....go around the unit circle \[5\pi \] times to get (-1,0)